OpenStudy (anonymous):

The width of a rectangle is 2 less than its length. If the perimeter of the rectangle is no more than30 centimeters, find the range of possible lengths and widths of the rectangle. Use the formula, P=2L+2W, for the perimeter of the rectangle. In the form of a paragraph, explain in complete sentences the steps necessary to set up the equation. Include the final inequality in your explanation.

OpenStudy (anonymous):

P = 2L + 2W 48 = 2L + 2W 48 = 2(12) + 2(12) 48 = 24 + 24 P = 48

OpenStudy (anonymous):

Next question plz

OpenStudy (anonymous):

stop hurrying people

OpenStudy (anonymous):

Thats not your business.. .;)

OpenStudy (anonymous):

=)